OpenStudy (anonymous):

NEED HELP FAST. -3x+2y=-15 2x-4y=26 elimination using multiplacation .

OpenStudy (tkhunny):

Multiply the first equation by 2 and look around for some magic.

OpenStudy (anonymous):

thank. you.

OpenStudy (anonymous):

4x=4/4=1. x=1

OpenStudy (tkhunny):

Okay, now do it a different way. Divide the second equation by 2.

OpenStudy (anonymous):

x-2y=13

OpenStudy (anonymous):

2y=11

OpenStudy (tkhunny):

And adding the two equations produces the same result as before?

OpenStudy (tkhunny):

Try to remember that there may be more than one way to proceed. Don't get all excited when you see a potential solution. Keep your eyes open for other ways to go about it.

OpenStudy (anonymous):

oh. -3(1)+2y=-15

OpenStudy (anonymous):

y=-4

OpenStudy (tkhunny):

Try out your complete solution in both ORIGINAL equations. That will tell you if you wandered off or if you nailed it.